OpenStudy (anonymous):

The length of a rectangle in inches is 16 less than 7 times the width. If the perimeter of the rectangle is 48 inches , what is the length in inches. A) 5 B) 8 C) 19 D) 40

OpenStudy (anonymous):

C p=2(l+w) 48=2(19+5)

OpenStudy (anonymous):

medal?